What is the perimeter of a rectangle if the length is 7/10 in. and the width is 1/10 in. 4/5, 1 4/5, 1 3/5, 3/5
Answer by ewatrrr(24785)


Hi,
*Note: the Perimeter of a rectangle is the sum of all of it's sides
7/10 + 7/10 + 1/10 + 1/10 = 16/10 = 8/5 = 1 3/5